OpenStudy (anonymous):

Assume that y varies inversely with x. If y = 7 when x = -3, find y when x = 7.

OpenStudy (jlastino):

First find the constant y=k/x (this how you write y varies inversely with x, with k being a constant) so 7=k/-3 -21=k y=-21/7 What is y? :D

OpenStudy (anonymous):

-3? :D

OpenStudy (jlastino):

Yeah :D